OpenStudy (anonymous):

write an equation in the form y=mx+b for the line that passes through the point (-5-1) and has a slope of -2

OpenStudy (anonymous):

ever heard of point-slope form

OpenStudy (anonymous):

y - y1=m(x-x1) where m= slope

OpenStudy (anonymous):

im in summer school right now and just need the answer please just help me out ha

OpenStudy (imer):

The site is called "openstudy" and not "openanswers" :P